# SSTP: Efficient Sample Selection for Trajectory Prediction

### v2

@@ 4 Experiments
Methods Ratio(%) HiVT-64 HiVT-128 HPNet
minADE minFDE xxMRxx minADE minFDE xxMRxx minADE minFDE xxMRxx
@@
SSTP (Ours) 0.711 1.088 0.114 0.696 1.048 0.106 0.671 0.931 0.076
Table 2: Performance comparison results on Argoverse 1 [5] with data retention ratios of 60%, 50%, and 40%. The compared methods include Random Selection [36], K-Means Clustering [24], and Herding Selection [2, 8]. The model used for data selection is HiVT-64 [61], while the evaluation is conducted on HiVT-64, HiVT-128, and HPNet [42]. Ratio (%) represents the proportion of retained data relative to the full training set.
Method Ratio(%) QCNet DeMo
minADE minFDE xxMRxx minADE minFDE xxMRxx
@@
Cluster 50 0.798 1.435 0.193 0.732 1.437 0.191
SSTP (Ours) 0.754 1.352 0.172 0.704 1.414 0.173
Random 0.811 1.471 0.226 0.763 1.475 0.202
Cluster 40 0.813 1.495 0.214 0.732 1.456 0.195
SSTP (Ours) 0.778 1.410 0.183 0.723 1.450 0.191
@@
Table 3: Performance comparison results on Argoverse 2 [47] with data retention ratios of 60%, 50%, and 40%. The compared methods include Random Selection and K-means clustering. The model used for data selection is QCNet [62], while the evaluation is conducted on QCNet and DeMo [50].
4.1 Benchmarks and Setup
Datasets. We evaluated the effectiveness of our proposed SSTP method on Argoverse Motion Forecasting Dataset 1.1  [5] and Argoverse 2 [47]. The Argoverse 1 dataset contains 323,557 real-world driving scenarios. All the training and validation scenarios are 5-second sequences sampled at 10 Hz. The length of the historical trajectory for each scenario is 2 seconds, and the length of the predicted future trajectory is 3 seconds. The Argoverse 2 dataset contains 250,000 scenarios, with the same sampling frequency of 10 Hz. Each trajectory has a larger observation window with 5 seconds and a longer prediction horizon with 6 seconds.
Baselines. For Argoverse 1, we validate our SSTP method on two SOTA models HiVT [61] and HPNet [42] for evaluation. For Argoverse 2, we evaluate our SSTP method using two SOTA models QCNet [62] and DeMo [50]. For a more comprehensive comparison, we also include three following data selection approaches:
(1) Random Selection [36]: randomly selects a certain proportion of training samples from the original training set.
(2) K-Means Clustering [24]: clusters trajectories within the observation window based on their features, and then selects the trajectory sample closest to the cluster center in each cluster as a representative.
(3) Herding Selection [2]: a greedy strategy that first computes the mean feature of all trajectories within the observation window and then iteratively selects trajectory samples that bring the mean of the selected subset as close as possible to the overall mean.
Implementation Details. We primarily utilized the pre-trained HiVT-64 and QCNet as backbone models to perform sample selection on the Argoverse 1 and Argoverse 2 datasets, respectively. To evaluate the performance of the selected subset, we follow their official training and validation protocols. We experimented with different selection ratios, different intervals, and assessed the prediction accuracy of the trajectory models after training on the corresponding subsets.
4.2 Main Results
Table 2 showcases the strong performance of our selected subset on the Argoverse 1 dataset across all compression rates. Following the same experimental setup as our baseline models, we trained HiVT and HPNet from scratch on the subset. When trained on the full dataset, HiVT-64 achieves a minADE of 0.695, minFDE of 1.037, and MR of 0.109. With a data retention rate of 60%, our selected subset achieves a minADE of 0.702, minFDE of 1.064, and MR of 0.110. This demonstrates that our method significantly reduces data volume while maintaining nearly lossless model performance. Even at a data retention rate of only 50%, models trained on our selected subset still maintain excellent performance, with minADE at 0.704, minFDE at 1.073, and MR at 0.111. Compared to random selection, our method reduces minADE by 0.046, minFDE by 0.102, and MR by 0.026. Moreover, while k-means clustering and herding methods show slight improvements in model performance, they still fall short of the performance achieved by our method. Furthermore, our subset also demonstrates superior performance on the HiVT128 model and the current SOTA model HPNet.
@@ 4.2 Main Results
We further evaluated our proposed method on the Argoverse 2 dataset, which presents greater challenges due to its more diverse driving scenarios and longer prediction horizons. As shown in Table 3, our method consistently outperforms other data selection strategies across all data retention rates, achieving lower minADE and minFDE while maintaining a lower MR. These results further validate the robustness of our approach, as it maintains strong performance across different datasets. This demonstrates that our method is not only effective within a specific dataset but also generalizes well to more complex and diverse trajectory scenarios, such as those found in Argoverse 2.
Agent$<$40 Agent>=>=> =40 Agent>=>=> =60 Agent>=>=> =80
minADE minFDE xMRx minADE minFDE xMRx minADE minFDE xMRx minADE minFDE xMRx
SSTP (50%) 0.716 1.102 0.111 0.962 1.497 0.183 1.219 1.835 0.280 1.373 1.762 0.277
Table 4: Comparison of model performance across different scene densities when trained on the full dataset, random selection for 50% subset versus the 50% subset selected by our method (SSTP), where ours achieves superior performance.
Variants Selection Strategy Data Distribution(%) Model Performance
Partition Submodular Agent$<$40 Agent>=>=> =40 minADE minFDE xMRx
Full dataset - - 93.88 6.12 0.692 1.047 0.104
Random - - 85.16 14.84 0.741 1.164 0.125
@@
Table 5: Performance comparison of different data selection strategies on HiVT trained with Argoverse 1. This table illustrates the impact of Partition and selection with Submodular Gain strategies on data distribution and model performance. The whole dataset and random selection serve as baselines, while different variations of SSTP are evaluated. Our method (SSTP), which integrates both strategies, achieves the best results by maintaining a balanced data distribution and reducing minADE, minFDE, and MR.
4.3 Ablation Study
Performance Enhancement. Scene density in autonomous driving varies significantly. However, most existing trajectory prediction datasets predominantly focus on low-density scenarios. From the perspective of safe driving, an ideal trajectory predictor should maintain strong performance across diverse driving scenarios, regardless of data density. To address this, we conducted a comprehensive evaluation of our proposed method on existing models and across different scene densities. As presented in Table 4, our method consistently outperforms models trained on the full dataset, particularly in high-density agent scenarios, across both Argoverse 1 and Argoverse 2 datasets. When the agent density is below 40, our method achieves comparable performance to models trained on the full dataset, with only marginal increases of 0.01 in minADE and 0.03 in minFDE, while MR remains nearly unchanged. However, in high-density scenarios where the agent count exceeds 60, models trained on our selected subset demonstrate significant improvements, with minADE and minFDE errors decreasing by over 0.03 and 0.06, respectively. This advantage becomes even more pronounced in scenarios with more than 80 agents, where minADE is reduced by approximately 0.08, minFDE by nearly 0.3, and MR by nearly 9%. Similarly, the SOTA model HPNet, when trained on our selected subset, outperforms its counterpart trained on the full dataset, demonstrating consistent improvements on the Argoverse 2 dataset.
@@ 4.3 Ablation Study
Density Balancing. Our method explicitly controls scene density distribution during selection, ensuring a more balanced dataset, as illustrated in Figure 2. In contrast, random selection fails to maintain this balance, resulting in an uneven distribution of scenarios with varying complexity, ultimately affecting the model’s generalization capability. As shown in Table 5 line 4, applying scene balancing alone already outperforms the random selection method. For instance, minADE decreases from 0.741 to 0.729, and minFDE drops from 1.164 to 1.116. This demonstrates that controlling scene density during data selection can enhance the effectiveness of trajectory prediction models. However, our full method outperforms scene balancing. By integrating submodular selection in addition to scene balancing, our approach further optimizes data selection by ensuring that the chosen samples carry higher information content. As a result, our method achieves minADE of 0.704 and minFDE of 1.073, further reducing errors. These findings indicate that while scene balancing is beneficial, it is insufficient to achieve optimal performance without also considering sample informativeness.

### v3

@@ 3 Experiments
3.1 Benchmarks and Setup
Datasets. We evaluated the effectiveness of our proposed SSTP method on Argoverse Motion Forecasting Dataset 1.1  (Chang et al., 2019) and Argoverse 2 (Wilson et al., 2023). The Argoverse 1 dataset contains 323,557 real-world driving scenarios. All the training and validation scenarios are 5-second sequences sampled at 10 Hz. The length of the historical trajectory for each scenario is 2 seconds, and the length of the predicted future trajectory is 3 seconds. The Argoverse 2 dataset contains 250,000 scenarios, with the same sampling frequency of 10 Hz. Each trajectory has a larger observation window with 5 seconds and a longer prediction horizon with 6 seconds.
Baselines. For Argoverse 1, we validate our SSTP method on two SOTA models HiVT (Zhou et al., 2022c) and HPNet (Tang et al., 2024) for evaluation. For Argoverse 2, we evaluate our SSTP method using two SOTA models QCNet (Zhou et al., 2023) and DeMo (Zhang et al., 2025). For a more comprehensive comparison, we also include three following data selection approaches:
(1) Random Selection (Rebuffi et al., 2017): randomly selects a certain proportion of training samples from the original training set.
(2) K-Means Clustering (Likas et al., 2003): clusters trajectories within the observation window based on their features, and then selects the trajectory sample closest to the cluster center in each cluster as a representative.
(3) Herding Selection (Castro et al., 2018): a greedy strategy that first computes the mean feature of all trajectories within the observation window and then iteratively selects trajectory samples that bring the mean of the selected subset as close as possible to the overall mean.
Methods $\alpha$(%) HiVT-64 HiVT-128 HPNet
[…]
@@ 3.1 Benchmarks and Setup
SSTP 0.711 1.088 0.114 0.696 1.048 0.106 0.671 0.931 0.076
Table 2: Performance comparison results on Argoverse 1 with data retention ratios of 60%, 50%, and 40%. The compared methods include Random Selection, K-Means Clustering, and Herding Selection. The model used for data selection is HiVT-64, while the evaluation is conducted on HiVT-64, HiVT-128, and HPNet. $\alpha$ (%) represents the proportion of retained data relative to the full training set.
Agent<<40 Agent>=>=40 Agent>=>=60 Agent>=>=80
[…]
Cluster 0.719 1.121 0.115 0.971 1.501 0.188 1.251 1.898 0.304 1.568 2.257 0.376
Herding 0.726 1.122 0.116 0.980 1.528 0.190 1.258 1.922 0.310 1.581 2.381 0.372
SSTP 0.714 1.098 0.113 0.962 1.497 0.183 1.219 1.835 0.280 1.373 1.762 0.277
Table 3: Comparison of model performance across scenario densities when trained on the full dataset, a 50% random subset, and a 50% SSTP subset, where SSTP consistently achieves superior results.
Implementation Details. We primarily utilized the pre-trained HiVT-64 and QCNet as backbone models to perform sample selection on the Argoverse 1 and Argoverse 2 datasets, respectively. To evaluate the performance of the selected subset, we follow their official training and validation protocols. We experimented with different selection ratios, different intervals, and assessed the prediction accuracy of the trajectory models after training on the corresponding subsets. Further implementation details are provided in Appendix A.1.
3.2 Main Results
We present the primary experimental results in this section, while a comprehensive ablation study, including analyses of alternative strategies and design choices, is provided in Appendix A.3. In addition, qualitative results and visual analyses are presented in Appendix A.4.
Table 2 showcases the strong performance of our selected subset on the Argoverse 1 dataset across all compression rates. Following the same experimental setup as the baseline models, we trained HiVT and HPNet from scratch on the subset. Our method significantly reduces data volume while maintaining nearly lossless performance. Even with only half of the data, models trained on our selected subset still perform comparably to those trained on the full dataset and consistently outperform random selection, clustering, and herding. Furthermore, our subset also demonstrates strong results on larger models such as HiVT-128 and HPNet. Additional results for other data retention ratios are provided in Appendix A.2.1, with detailed numbers reported in Table 6.
Method $\alpha$(%) QCNet DeMo
[…]
@@ 3.2 Main Results
Cluster 0.798 1.435 0.193 0.732 1.437 0.191
Herding 0.782 1.407 0.190 0.731 1.434 0.190
SSTP 0.754 1.352 0.172 0.704 1.414 0.173
Random 40 0.811 1.471 0.226 0.763 1.475 0.202
Cluster 0.813 1.495 0.214 0.732 1.456 0.195
Herding 0.807 1.454 0.198 0.739 1.460 0.197
SSTP 0.778 1.410 0.183 0.723 1.450 0.191
@@
Table 4: Performance comparison results on Argoverse 2 with data retention ratios of 60%, 50%, and 40%. The compared methods include Random Selection and K-means clustering. The model used for data selection is QCNet, while the evaluation is conducted on QCNet and DeMo.
Performance Enhancement. Scene density in autonomous driving varies substantially, yet most existing trajectory prediction datasets are dominated by low-density scenarios. From a safety perspective, however, an ideal trajectory predictor must perform reliably across the full spectrum of scene complexities. To this end, we evaluate our proposed method on multiple models and across different density levels. As shown in Table 3, our method consistently outperforms models trained on the full dataset, particularly in high-density scenarios, on both the Argoverse 1 and Argoverse 2 datasets. In low-density settings (fewer than 40 agents), our selected subset achieves performance nearly identical to the full dataset, with only marginal increases of minADE and minFDE, while MR remains almost unchanged, which is negligible given the limited interactions in such scenes.
When the agent density increases, our method brings the most substantial gains. Compared to all other baselines, SSTP achieves lower displacement errors and notably reduces MR. For example, when the number of agents exceeds 80, SSTP cuts the missing rate by more than 8% relative to random selection, and also outperforms clustering and herding by clear margins. These results highlight that while clustering and herding provide partial improvements by ensuring representativeness, they are still insufficient for handling highly complex traffic scenes. In contrast, SSTP effectively balances scene density while selecting informative samples, leading to consistently superior performance across all density levels. Similarly, the SOTA model HPNet, when trained on our selected subset, even outperforms its counterpart trained on the full dataset. Detailed results are provided in Appendix A.2.2, with comprehensive numbers reported in Table 8.
Figure 3: Performance of HiVT and HPNet trained on subsets selected by different methods at varying retention ratios, evaluated with minADE.
Figure 4: Distribution of scenarios categorized by agent density in Argoverse 1 before (left) and after (right) applying SSTP with 50% data retention ratio.
@@
Generalization Across Datasets. We further evaluated our proposed method on the Argoverse 2 dataset, which presents greater challenges due to its more diverse driving scenarios and longer prediction horizons. As shown in Table 4, our method consistently outperforms other data selection strategies across all data retention rates, achieving lower minADE and minFDE while maintaining a lower MR. These results further validate the robustness of our approach, as it maintains strong performance across different datasets. This demonstrates that our method is not only effective within a specific dataset but also generalizes well to more complex and diverse trajectory scenarios, such as those found in Argoverse 2. Additional results for other data retention ratios are provided in Appendix A.2.1, with detailed numbers reported in Table 7.
Variants Data Distribution(%) Model Performance
[…]
Full dataset 93.88 6.12 0.692 1.047 0.104
Random 85.16 14.84 0.741 1.164 0.125
@@
Table 5: Performance comparison of data selection strategies on HiVT trained with Argoverse 1. This table shows the impact of partitioning and selection with Submodular Gain strategies on data distribution and model performance. The whole dataset and random selection serve as baselines, while different variations of SSTP are evaluated. Our method (SSTP), integrates both strategies, achieves the best results by maintaining a balanced data distribution and reducing minADE, minFDE, and MR.
Data Retention Ratio. To examine the impact of different data retention ratios on model performance, we conducted experiments with retention rates of $\alpha$ = 60, 50, 40, 30, 20, 10% as shown in Figure 4. When higher model performance is required, retaining 50% of the data already achieves results comparable to training with the full dataset. This demonstrates the effectiveness of our SSTP method, as the selected subset is of higher quality compared to equally sized subsets chosen by other methods. Furthermore, under limited computational resources, even retaining only 20% of the data still yields reasonably good results. We further analyze the behavior under extremely low retention ratios, with detailed results provided in Appendix A.2.3 and Table 9.
@@
Density Balancing. Our method explicitly controls scene density distribution during selection, ensuring a more balanced dataset, as illustrated in Figure 4. In contrast, random selection fails to maintain this balance, leading to uneven representation of scenarios with varying complexity and weaker generalization. As shown in Table 5 line 4, applying scene balancing alone already improves performance compared to random selection, demonstrating that controlling scene density enhances the effectiveness of trajectory prediction models. However, balancing alone remains insufficient. By further integrating submodular selection to account for sample informativeness, our method achieves the best performance across all metrics. These findings indicate that while scene balancing is beneficial, it is insufficient to achieve optimal performance without also considering sample informativeness.
Figure 5: Training time comparison of HiVT and HPNet. Our method achieves over 50% reduction in time while maintaining comparable accuracy.
Efficiency. Our method significantly reduces computational time while maintaining strong performance, as shown in Figure 5. Training the HiVT model on the full dataset requires 7.78 hours. In contrast, utilizing our SSTP method to select a 60% subset requires only 0.6 hour for pre-training and 2.30 hours for selection, reducing the overall training time. When training on the selected subset, the total training time decreases to 7.25 hours (0.6 + 2.30 + 4.35), with only a minor increase of 0.007 in minADE.
Beyond efficiency on the original backbone, a key advantage of SSTP is that the selected subset is model-agnostic and can be directly reused to train other trajectory predictors. For instance, when applied to HPNet, training on the full dataset takes 307.2 hours, whereas training with the selected subset reduces the time by over 100 hours, cutting the training cost by nearly 45%, while achieving comparable or even better accuracy. This decisively confirms the superior efficiency of our approach in balancing training cost and model performance.
Generalizability across Backbones. To further assess the generalizability of our method, we evaluated its performance with different backbone models. Specifically, we used pre-trained HiVT-64 and HPNet as feature extractors on the Argoverse 1 dataset and conducted experiments under varying data retention ratios. Detailed results are provided in Appendix A.2.4, with numerical values reported in Table 10. As shown, regardless of the backbone used for subset selection, the final trajectory prediction performance remains nearly identical. This demonstrates that our subset selection strategy is largely backbone-agnostic, underscoring its robustness and broad applicability for optimizing diverse trajectory prediction models.
4 Related Work
Trajectory Prediction. Trajectory prediction infers an agent’s future motion from its historical observations. In recent years, research has increasingly concentrated on capturing complex multi-agent interactions, driving advances in various predictive methods (Wang et al., 2024; Giuliari et al., 2021; Xu et al., 2020; Liu et al., 2021; Ngiam et al., 2021; Zhou et al., 2022c; 2023; Tang et al., 2024; Zhang et al., 2025; Liang et al., 2020). Furthermore, novel approaches including pretraining (Chen et al., 2023; Cheng et al., 2023; Lan et al., 2023), historical prediction structure design (Park et al., 2024; Tang et al., 2024), GPT-style next-token prediction (Philion et al., 2023; Seff et al., 2023), and post-processing optimization (Zhou et al., 2024; Choi et al., 2023) have significantly enhanced model performance, demonstrating strong results across various datasets. However, most state-of-the-art methods rely on training with large-scale datasets (Caesar et al., 2020; Chang et al., 2019; Sun et al., 2020; Wilson et al., 2023; Ettinger et al., 2021), leading to significant computational costs. In contrast, we propose a pioneering data selection strategy in the trajectory prediction domain that constructs a compact, balanced, yet highly representative dataset, significantly reducing training time while preserving model performance.
Long-Tail in Trajectory Prediction Dataset The performance of the trajectory prediction models is evaluated on the overall average. While they excel on benchmarks, these models often struggle with challenging scenarios (Makansi et al., 2021; Wang et al., 2023; Pourkeshavarz et al., 2023) due to the long-tail data distribution, where common cases dominate and complex or rare situations are underrepresented (Chen et al., 2024). Recent studies (Wang et al., 2023; Zhang et al., 2024; Zhou et al., 2022a; Lan et al., 2024) have started addressing the long-tail problem in trajectory prediction, primarily by leveraging contrastive learning to enhance feature representations. However, these methods mainly focus on optimizing feature-level learning while overlooking scenario level distribution and the importance of individual samples. In contrast, our proposed method assesses the contribution of each sample and applies a refined selection strategy to build a balanced and compact dataset. Experimental results show that it significantly boosts performance in complex scenarios.
Training Sample Selection Deep neural networks, especially Transformers, depend on large datasets and incur high computational costs. To reduce these costs and shorten training time, various methods to improve data efficiency have been proposed, including frequent parameter updates (Robbins & Monro, 1951), fewer iterations (Sutskever et al., 2013), and dynamic learning rate adjustments (Kingma, 2014; Duchi et al., 2011). To directly reduce data volume, dataset condensation compresses raw data into compact synthetic samples (Wang et al., 2018; Zhao et al., 2020; Zhao & Bilen, 2021; Kim et al., 2022; Wang et al., 2022; Zhao & Bilen, 2023; Cazenavette et al., 2022). Another widely studied approach is coreset selection, which constructs a weighted subset that closely approximates the statistical distribution of the original dataset (Har-Peled & Mazumdar, 2004; Coleman et al., 2019; Margatina et al., 2021; Mirzasoleiman et al., 2020). However, those approaches on dataset condensation (Nguyen et al., 2020; Loo et al., 2022; Zhou et al., 2022b; Sajedi et al., 2023; Zhao et al., 2023) and coreset selection (Killamsetty et al., 2021; Paul et al., 2021) has been predominantly focused on image classification. In contrast, we propose a sample selection strategy based on submodular functions for the trajectory prediction domain. As a pioneering method, our approach significantly reduces the training data required while maintaining model performance.
5 Conclusion
Limitations. While effective, our method still incurs non-negligible computational overhead during the gradient extraction and submodular optimization stages of sample selection. Although this is a one-time cost, further optimization to reduce its complexity would be desirable. Moreover, model performance degrades under extremely low data retention ratios (e.g., 10%), highlighting the challenge of preserving robustness when only very limited data are available. Finally, the present study is restricted to trajectory prediction tasks on the Argoverse benchmarks; extending SSTP to other autonomous driving tasks and additional datasets remains an important avenue for future research.
Ethics Statement
This work uses only publicly available autonomous driving datasets (Argoverse 1 and 2) containing vehicle trajectories and maps, with no human subjects or private data involved. We adhere to the conference Code of Ethics and confirm full compliance with its standards.
Reproducibility Statement
Implementation details, including dataset preprocessing, training protocols, and hyperparameter settings, are provided in Section 4 and Appendix A.1. We also release our code and checkpoints (via the anonymous link) to facilitate full reproducibility.
Appendix A Appendix
A.1 Implementation Details
During the dataset selection stage, we employ HiVT (Zhou et al., 2022c), HPNet (Tang et al., 2024), and QCNet (Zhou et al., 2023) as backbones with batch sizes of 16, 4, and 4, respectively, selected using the NaiveGreedy algorithm. For model evaluation, one mainstream model and one SOTA model are chosen per dataset: HiVT and HPNet for Argoverse 1 (Chang et al., 2019), and QCNet and DeMo (Zhang et al., 2025) for Argoverse 2 (Wilson et al., 2023). All models follow their official experimental settings and are trained on both selected subsets and complete datasets.
HiVT: Batch size 32, LR $3\times 10^{-4}$, weight decay $1\times 10^{-4}$, dropout 0.1. Its architecture includes 1 interaction module, 4 temporal and 3 global modules, 8-head attention (50 m radius), 6 prediction modes, and hidden sizes of 64/128.
HPNet: Batch size 16, LR $5\times 10^{-4}$ (same weight decay and dropout). It has 1 spatiotemporal and 2 tri-factor attention layers (50 m radius, 20-frame windows) with data augmentation: horizontal flipping (0.5), agent (0.05) and lane (0.2) occlusion.
QCNet: Trained for 64 epochs with AdamW (batch size 32, LR $5\times 10^{-4}$, weight decay $1\times 10^{-4}$, dropout 0.1). It uses a 128-dim hidden space, gated 8-head attention, layer normalization on MLP and attention layers, 3 iterations for the trajectory proposal, and 2 multi-context attention modules in both encoder and decoder.
DeMo: Trained for 60 epochs with AdamW (batch size 16 per GPU, LR 0.003, weight decay 0.01, cosine annealing with 10 warm-up epochs). No data augmentation is applied.
A.2 Additional Experiments
A.2.1 Additional Results on Argoverse 1 and Argoverse 2.
We present comprehensive results for different models and data retention ratios on SSTP-selected datasets for both Argoverse 1 and Argoverse 2. The results, shown in Table 6 and Table 7, follow trends consistent with those observed in the 60%, 50%, and 40% retention settings, further validating the effectiveness of our approach.
Methods $\alpha$(%) HiVT-64 HiVT-128 HPNet
[…]
Random 30 0.76 1.21 0.13 0.73 1.13 0.12 0.69 0.98 0.09
Cluster 0.74 1.14 0.12 0.71 1.07 0.11 0.68 0.97 0.08
Herding 0.73 1.14 0.12 0.71 1.07 0.11 0.69 0.96 0.09
SSTP 0.72 1.12 0.12 0.70 1.06 0.11 0.68 0.95 0.07
Random 20 0.78 1.25 0.14 0.75 1.18 0.13 0.71 1.01 0.10
Cluster 0.76 1.21 0.14 0.72 1.11 0.12 0.69 0.98 0.09
Herding 0.76 1.19 0.13 0.74 1.14 0.13 0.70 0.99 0.09
SSTP 0.74 1.16 0.12 0.72 1.11 0.11 0.69 0.97 0.08
Random 10 0.84 1.40 0.16 0.80 1.29 0.15 0.88 1.40 0.18
Cluster 0.82 1.37 0.16 0.76 1.20 0.13 0.87 1.39 0.17
Herding 0.80 1.29 0.14 0.80 1.29 0.14 0.81 1.26 0.13
SSTP 0.78 1.25 0.13 0.75 1.17 0.12 0.78 1.17 0.12
Table 6: Performance comparison results on Argoverse 1 with data retention ratios of 30%, 20%, and 10%. Pretrained HiVT-64 is used for sample selection. Evaluation conducted on HiVT-64, HiVT-128, and HPNet.
A.2.2 Performance Enhancement on HPNet.
To further validate the generalization capability of the SSTP method, we conducted the same experimental setup on HPNet. The test set was partitioned based on different scene densities, and the model was evaluated on these subdivided datasets. The results, presented in Table 8, demonstrate the effectiveness of our approach. When the agent density is below 40, our method achieves performance comparable to models trained on the full dataset, with only slight increases of 0.019 in minADE and 0.028 in minFDE, while MR remains nearly unchanged. However, in high-density scenarios where the number of agents exceeds 60, models trained on our selected subset exhibit notable improvements, with minADE and minFDE reduced by 0.001 and 0.002, respectively. This advantage becomes even more pronounced in scenarios with more than 80 agents, where minADE is reduced by approximately 0.022, minFDE by nearly 0.073, and MR by nearly 3.5%. These findings confirm that SSTP effectively maintains dataset diversity while preserving model performance, regardless of the model used.
Methods $\alpha$(%) QCNet DeMo
[…]
Random 30 0.827 1.486 0.224 0.779 1.547 0.216
Cluster 0.821 1.523 0.222 0.765 1.498 0.207
Herding 0.819 1.505 0.213 0.762 1.506 0.204
SSTP 0.794 1.453 0.192 0.743 1.501 0.192
Random 20 0.878 1.586 0.242 0.827 1.645 0.238
Cluster 0.843 1.561 0.222 0.794 1.618 0.229
Herding 0.844 1.545 0.222 0.801 1.625 0.227
SSTP 0.832 1.530 0.211 0.783 1.591 0.212
Random 10 0.944 1.797 0.272 0.918 1.817 0.263
Cluster 0.911 1.772 0.251 0.878 1.779 0.247
Herding 0.908 1.743 0.246 0.882 1.774 0.243
SSTP 0.891 1.684 0.241 0.871 1.770 0.241
Table 7: Performance comparison results on Argoverse 2 with data retention ratios of 30%, 20%, and 10%. Pretrained QCNet is used for sample selection. Evaluation conducted on QCNet and DeMo.
HPNet Agent<<40 Agent>=>=40 Agent>=>=60 Agent>=>=80
[…]
Full 0.611 0.833 0.062 0.875 1.202 0.127 1.111 1.463 0.201 1.596 1.771 0.276
Random 0.652 0.903 0.071 0.932 1.336 0.147 1.267 1.850 0.274 1.649 2.086 0.375
Cluster 0.642 0.889 0.069 0.916 1.291 0.143 1.193 1.669 0.250 1.636 1.977 0.325
Herding 0.650 0.901 0.070 0.921 1.303 0.145 1.215 1.673 0.252 1.637 2.001 0.342
SSTP 0.630 0.861 0.064 0.878 1.210 0.132 1.110 1.461 0.198 1.574 1.698 0.241
Table 8: Performance Comparison of HPNet across different scene densities when trained on different sample set with $\alpha=50\%$.
[…]
10 Random 0.818 1.329 0.167
SSTP 0.781 1.253 0.132
5 Random 0.869 1.484 0.183
SSTP 0.843 1.416 0.161
2 Random 1.055 1.902 0.245
SSTP 1.015 1.871 0.220
1 Random 2.033 4.613 0.452
SSTP 1.589 3.416 0.386
Table 9: Performance comparison under extremely low data retention settings ($\alpha\leq 10\%$). For each retention ratio, SSTP consistently outperforms random selection, though performance degrades sharply when the data retention ratio drops below 10%.
Backbone $\alpha$(%) HiVT-64 HPNet
[…]
HiVT-64 10 0.786 1.254 0.138 0.782 1.172 0.121
20 0.749 1.167 0.125 0.691 0.970 0.081
30 0.727 1.120 0.120 0.678 0.951 0.079
40 0.712 1.089 0.114 0.671 0.931 0.076
50 0.704 1.073 0.111 0.661 0.913 0.074
60 0.703 1.065 0.111 0.654 0.901 0.072
HPNet 10 0.796 1.277 0.143 0.739 1.050 0.095
20 0.753 1.193 0.125 0.699 0.975 0.082
30 0.732 1.143 0.119 0.681 0.943 0.078
40 0.717 1.087 0.112 0.670 0.924 0.075
50 0.708 1.079 0.111 0.664 0.917 0.074
60 0.703 1.067 0.108 0.657 0.910 0.075
Table 10: Performance comparison of different backbone models, HiVT-64 and HPNet, trained on subsets selected by different strategies at varying data retention ratios.
A.2.3 Low Data Retention Setting.
We further studied the extremely low-budget regime with $\alpha\in\{10,5,2,1\}\%$. As shown in Table 9, the performance of all methods drops sharply below 10%. Nevertheless, our SSTP method consistently outperforms random selection at the same budget levels. In particular, at 1% retention, SSTP reduces minADE from 2.033 to 1.589 and MR from 0.452 to 0.386, showing that even with very limited data, informative subset selection remains beneficial. This advantage comes from prioritizing scarce high-density scenes during budget allocation, ensuring long-tailed scenarios remain represented.
A.2.4 Generalizability and Robustness.
To further assess the generalizability and robustness of our proposed method, we evaluate the impact of different data retention ratios on SSTP-selected datasets using various backbone networks. The results, presented in Table 10, demonstrate that SSTP is not only adaptable across a diverse range of trajectory prediction models but also effectively reduces dataset size while maintaining model performance across different feature extractors. These findings underscore SSTP’s robustness and broad applicability in real-world autonomous driving scenarios.
A.3 Ablation Study